# Brain-inspired Large Language Model

### v6

@@ 5 Conclusion, Limitation and the Future
BriLLM introduces a transformative framework for language modeling, replacing attention-based architectures with brain-inspired dynamic signal propagation over a fully connected graph. By representing tokens as nodes and using energy dynamics to guide signal flow, BriLLM enables non-autoregressive generation, full node-level interpretability, and theoretically infinite $n$-gram modeling. Its design decouples model size from sequence length, optimizing resource use while mimicking neurocognitive processes like memory formation. This work challenges the dominance of attention mechanisms, offering a scalable, transparent alternative aligned with neural signaling principles.
Current limitations stem from constrained computational resources, with early checkpoints using moderate hyperparameters—yet performance already matches GPT-1 (Radford et al., 2018). A key distinction from traditional machine learning lies in SiFu’s support for concurrent multi-input/multi-output streams, whereas conventional models process a single input-output pair at a time. We envision an embodied BriLLM where multiple independent signal streams propagate simultaneously, guided by energy maximization—enabling parallel thought processes analogous to human multitasking, a capability beyond traditional single-stream architectures.
SiFu learning differs fundamentally from representation learning in deep learning. In BriLLM, the primary learnable parameter is the bias vector $b$, which lacks explicit representational intent; its role is to filter signal flows rather than encode semantic content. Even if interpreted as a weak form of node embedding, $b$ is far removed from the robust representational frameworks central to deep learning.
BriLLM’s current parameter count is ($n\times(1+d_{node}))^{2}+n^{2}\times d_{node}$ (with $n$ as token count and $d_{node}$ as node dimension), leading to quadratic growth. However, sparse training—using shared parameters for low-frequency token pairs—reduces size by  90% (Table 2), demonstrating potential for efficient scaling.
Training efficiency remains a challenge: BriLLM’s dynamic per-sample network construction hinders parallelization, and energy-based ranking over large vocabularies demands extensive computations. Approximate ranking—comparing candidate tokens only to relevant counterparts—could mitigate this, enabling parallel training of non-overlapping local networks.
Note that even though BriLLM theoretically supports infinite-gram language model without increasing model size, in practice, the model during training has to cover long enough input sequences, otherwise BriLLM decoding cannot give good enough sequence prediction beyond the training sample length. However, facilitating longer sequence prediction in terms of BriLLM simply depends on longer training without resizing the model itself.
BriLLM’s full interpretability enables innate multi-modal integration: nodes need not be limited to linguistic tokens but can represent units from diverse modalities. Critically, redefining nodes for new modalities does not require retraining from scratch—unlike LLMs, which rely on input/output alignment.
The initial signal vector (Eq. 1) is uniformly initialized, but we have validated stability with random initialization. Future work will explore signal shaping as a form of prompt engineering, potentially enabling more natural context specification than in-context learning in LLMs.
BriLLM’s performance at 1-2B parameters, comparable to 0.1B GPT-1, suggests a high threshold for LLM-like emergent capabilities. Additionally, supervised fine-tuning (SFT) methods for BriLLM remain underexplored, representing a key area for future research.
Table 5: Comparison of GPT-LLM and BriLLM.
LLM BriLLM
@@
model size correlated to input context length independent
interpretability only in input & output all nodes throughout the model
multi-modal implementation limited to be joined from input/output all nodes throughout the model

### v7

@@ 5 Conclusion, Limitation and the Future
BriLLM and its SiFu foundation represent a paradigmatic shift, addressing AGI’s core barrier: the black-box nature of conventional ML/DL. By rewriting learning around brain-inspired principles—static semantic nodes (like cortical regions Huth et al. (2016)) and dynamic signal propagation — we achieve full interpretability, unbounded context, and global-scale brain-like processing. This third innovation — modeling the brain as a system-level processor — addresses a critical gap in AI research, where prior work has not attempted to replicate the brain’s global operations.
BriLLM’s design directly mirrors two defining properties of the brain: static mapping of semantic units to distinct components (nodes, analogous to cortical regions Huth et al. (2016)) and dynamic signal propagation (analogous to electrophysiological activity) driving cognition. This enables three key capabilities absent in conventional LLMs: full interpretability across all components, decoupling of model size from sequence length, and inherent multi-modal compatibility (since nodes can represent any semantic unit, not just language).
Our initial 1–2B parameter models validate the design of BriLLM: they replicate GPT-1’s core generative capability (sequence continuation) with stable learning dynamics, despite being engineered to target GPT-3-level performance. Limitations reflect early-stage development (sub-scale size, sparse training refinement) rather than fundamental flaws. Additionally, while BriLLM theoretically handles infinite sequences, practical performance on very long sequences requires extended training on longer samples — consistent with the brain’s need for experience to develop long-term reasoning.
Future work will: (1) scale to larger checkpoints to test emergent abilities; (2) add multi-modal nodes for cross-modal processing; (3) refine signaling to mimic neural plasticity; and (4) develop embodied versions with sensorimotor integration.
Table 5 summarizes BriLLM’s advantages over conventional LLMs, highlighting its breakthrough in replicating the brain’s global properties. By redefining language modeling as a simulation of the brain’s macroscopic mechanisms, BriLLM paves the way for AGI rooted in the principles of biological intelligence.
Table 5: GPT-LLM vs. BriLLM comparison
GPT-LLM BriLLM
@@
Model size Tied to context length Context-independent (brain-like)
Interpretability Input/output only Full node-level transparency (cortical analogy)
Multi-modality Input/output aligned Native cross-modal nodes (brain integration)
Long sequences Quadratic complexity Linear complexity (signal overlay)
Error tracing Ambiguous (e.g., attention) Specific signal paths (neuroimage-like)